# NeSyGeo: A Neuro-Symbolic Framework for Multimodal Geometric Reasoning Data Generation

### v1

@@ Appendix B Statistics of NeSyGeo-Caption and NeSyGeo-CoT
Detailed numerical statistics and element distribution for the NeSyGeo-Caption and NeSyGeo-CoT datasets are presented in Table 6 and  6. For element distribution statistics, we randomly sampled 1.8k Geo-DSL sequences corresponding to images from each dataset, counting the frequency of different geometric elements. To facilitate interpretation, these elements are converted into corresponding natural language descriptions.
@@
Appendix C Additional Experimental Details and Results
We utilized the VLM-R1 [22] framework for RL experiments, conducted on 6 vGPU-32 GB. We set epochs to 2, num generations to 6, batchsize to 1. To enhance the visual perception capabilities of MLLMs, parameters of the language model and vision modules are set to be trainable.
@@ Appendix C Additional Experimental Details and Results
For SFT experiments, we employed the LLaMA-Factory [40] framework on 2 A800 GPUs with LoRA. We set the learning rate of $1\times 10^{-5}$, LoRA rank of 64, and use Adam optimization. Training on NeSyGeo-Caption used 1 epoch, while NeSyGeo-CoT used 2 epochs.
To investigate visual diversity directly across datasets, we randomly sampled 1k images from each, extracted features using ResNet, and visualized them via t-SNE. As illustrated in Figure 10, G-LLaVA—having augmented only the textual components of its base dataset—displays a distinctly non-uniform distribution in the image feature space. Conversely, our method exhibits uniform feature distributions, underscoring the substantial visual diversity of images generated by our approach.
@@
Figure 10: T-SNE of the image features of different automatic frameworks. Our datasets exhibit uniform feature distributions, underscoring the substantial visual diversity of images generated by the NeSyGeo framework.
Appendix D More Examples of NeSyGeo-CoT Dataset
Due to LLMs’ powerful search and reasoning capabilities, we obtain diverse Q&A pairs concerning properties like lengths, angles, or areas alongside high-quality CoT step-by-step. This process, involving backwards search across the geometric space defined by the symbolic form and forward validation, ensures the correctness of the numerical answers, thereby enriching textual diversity.
[…]
where perception difficulty is the number of Geo-DSL statements, and reasoning difficulty is the number of reasoning steps.
Appendix E Details of NeSyGeo-Test Benchmark.
Our NeSyGeo-Test benchmark comprises 2668 Q&A pairs. Consistent with the training set, numerical annotations are embedded in the image space, with only essential conditions and questions provided in the text. The type of numerical quantity categorizes the dataset sought: Angle (658 pairs), Shape (730 pairs), and Length (1280 pairs). Shape type includes shape area and perimeter, while length type includes edge and arc lengths. Based on the difficulty level in  D, problem difficulty is divided into three levels: Easy (1537 pairs), Medium (908 pairs), and Hard (223 pairs). Evaluation results on current mainstream open-source and closed-source MLLMs are shown in Table 8.
@@ Appendix D More Examples of NeSyGeo-CoT Dataset
Qwen2.5-VL-7B-Instruct 7B 38.3 30.7 32.2 36.8 27.5 32.3 43.3
InternVL2.5-4B 4B 53.8 53.1 53.5 62.9 49.4 32.0 53.4
InternVL2.5-8B 8B 55.9 56.2 55.4 64.5 49.4 43.3 55.8
LLaVA-NeXT-7B 7B 23.7 15.5 15.4 18.6 14.7 13.3 6.4
LLaVA-NeXT-13B 13B 15.7 15.6 16.0 15.8 16.0 15.2 15.8
@@
Appendix F Limitations
  • Limited Training Paradigm: Our current evaluation of the NeSyGeo dataset relies on a simple training paradigm to assess its efficacy for automated data generation. This approach lacks advanced training strategies, such as CLIP alignment or curriculum learning, which restricts the development of specialized models optimized for geometric reasoning tasks.
  • Restricted Domain Scope: The NeSyGeo framework is currently tailored to plane geometry, limiting its generalizability to other domains. However, we believe that for multimodal datasets in other domains, we can similarly achieve synthesis by defining symbolic statements, shifting the synthesis process to a controllable symbolic space, and constructing a symbolic-to-image engine, which we plan to explore in our future work.
  • Dependency on External APIs: The construction of Q&A pairs in this study partly relies on the reasoning capabilities of LLMs. This dependency increases generation costs and introduces potential inconsistencies. We aim to develop an automated solver that conducts search and validation directly within the symbolic space, thereby removing reliance on LLMs. This could further reduce costs and ensure complete rigor.
Appendix G Details of Prompts in Reverse Search and Forward Validation
In our automatic synthesis framework, we employ DeepSeek R1 as the expert LLM for reverse search and DeepSeek V3 for forward validation. The specific prompts utilized are detailed in Figures 12 and 13, respectively. Note that the blue text in these prompts is substituted with actual content.
Appendix H Detailed Definition of Geo-DSL
@@
Geo-DSL adopts an entity-relation-constraint framework to define geometric elements in plane geometry, encompassing 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes. Representative examples of symbolic statements and their corresponding natural language descriptions are illustrated in Figures 18, 15, 16, and 17. With a single statement, Geo-DSL uniquely specifies spatial elements, ensuring the accuracy of geometric synthesis while significantly facilitating parsing and transformation by our conversion engine. This language achieves comprehensive coverage of plane geometry, including numerical attributes such as lengths and angle measures, enabling precise and complete geometric representations. By streamlining definitions into concise statements, Geo-DSL reduces the complexity of symbolic processing, enhances the efficiency of the conversion engine, and supports seamless integration with neural synthesis pipelines. These advantages make Geo-DSL a robust and versatile solution for generating high-quality multimodal geometric reasoning data.
Appendix I Detailed Actions in Symbolic Spaces
@@
Figure 12: Prompt for Deepseek R1 in reverse search.
Figure 13: Prompt for Deepseek V3 in forward validation.

### v2

@@ B.1 Statistics of Datasets
Detailed numerical statistics and element distribution for the NeSyGeo-Caption and NeSyGeo-CoT datasets are presented in Table 7 and  7. For element distribution statistics, we randomly sampled 1.8k Geo-DSL sequences corresponding to images from each dataset, counting the frequency of different geometric elements. To facilitate interpretation, these elements are converted into corresponding natural language descriptions.
@@
B.2 Difficulty Definitions
$\alpha\times\text{perception difficulty}+(1-\alpha)\times\text{reasoning difficulty},$ (3)
where perception difficulty is the number of Geo-DSL statements, and reasoning difficulty is the number of reasoning steps. Here, we set alpha as 0.3.
B.3 Details of NeSyGeo-Test Benchmark.
Our NeSyGeo-Test benchmark comprises 2668 Q&A pairs. Consistent with the training set, numerical annotations are embedded in the image space, with only essential conditions and questions provided in the text. The type of numerical quantity categorizes the dataset sought: Angle (658 pairs), Shape (730 pairs), and Length (1280 pairs). Shape type includes shape area and perimeter, while length type includes edge and arc lengths. Based on the difficulty level in  B.2, problem difficulty is divided into three levels: Easy (1537 pairs), Medium (908 pairs), and Hard (223 pairs).
Appendix C More Details and Results of Experiments
C.1 Settings of Dataset Generation
To synthesize a diverse dataset, we set hyperparameters for generating images by configuring the step count, length range, and angle range. The step count $N$ ranges from one to four. The line length is defined within the basic range $l_{\text{min}},l_{\text{max}}]=[1,5]$, scalable by any multiple of 2 (e.g., $[4,20]$). The angle is constrained to multiples of 15° within $15^{\circ},165^{\circ}]$, with increased weights assigned to special angles. The temperature hyperparameters $\tau_{e}$ and $\tau_{a}$ were set to a default value of 1. We generate various types of weight matrices $A$ and $I$ by adjusting their corresponding values.
C.2 Human Preference evaluation
To rigorously assess the quality of our generated data, we conducted a comprehensive human evaluation study. We compared our NeSyGeo-CoT dataset against two prominent, automatically generated geometry datasets: G-LLaVA and MAVIS.
Methodology
We recruited 40 volunteers. To ensure a proficient and consistent level of domain expertise, a prerequisite for participation was holding a university undergraduate degree or higher. To ensure the objectivity and accuracy of the evaluation, participants were informed that the study pertained to middle school mathematics education and that their primary task was to meticulously verify the correctness of the provided solutions and final answers for each problem. We guaranteed participants that their responses would be completely anonymized to protect their privacy, with the right to withdraw at any time without penalty.
Each participant was presented with a questionnaire containing nine geometric problems. To mitigate potential ordering effects and sampling bias, the problems were composed of three samples randomly and independently drawn from each of the three datasets being evaluated.
Evaluation Criteria
Participants were asked to rate each sample on a 5-point Likert scale across five key dimensions designed to capture the multifaceted quality of a multimodal reasoning problem:
Image
This dimension evaluates the visual output. Image Correctness assesses whether the image 1)accurately reflects all entities and constraints from the text premise.2) has no conflict itself Image Appearance rates the human aesthetic quality, including the clarity of annotations and overall layout.
Question
This assesses the generated question. Question Correctness evaluates its logical validity and solvability based on the provided context, while Question Fluency rates its grammatical structure and naturalness.
Reasoning
This pertains to the generated solution. Reasoning Correctness scrutinizes the logical and mathematical validity of the step-by-step thought process. Reasoning Fluency assesses the clarity and readability of the explanation.
Overall
This captures the evaluators’ subjective impressions. Overall Reality gauges the plausibility of the problem, assessing if it feels like a genuine question from a textbook or exam. Overall Difficulty is the participant’s rating of the cognitive challenge involved.
Answer
Answer Correctness provides the definitive assessment of problem-solving capability by verifying if the final numerical answer derived from the reasoning process is objectively correct.
Results The aggregated results are presented in Picture 4. The findings indicate a clear preference for our NeSyGeo-CoT dataset. While all datasets scored perfectly on answer correctness—an expected outcome for synthetically generated data—our dataset was rated significantly higher in Reasoning Path Quality and Overall Difficulty. This suggests that our framework not only generates valid and high-fidelity problems but also produces content that is more complex and challenging, thereby validating its effectiveness for advancing the reasoning capabilities of MLLMs.
C.3 Reinforcement Learning
We utilized the VLM-R1 (Shen et al., 2025) framework for RL experiments, conducted on 6 vGPU-32 GB. We set epochs to 2, num generations to 6, batchsize to 1. To enhance the visual perception capabilities of MLLMs, parameters of the language model and vision modules are set to be trainable.
@@ C.3 Reinforcement Learning
C.4 Supervised Finetuning
For SFT experiments, we employed the LLaMA-Factory (Zheng et al., 2024) framework on 2 A800 GPUs with LoRA. We set the learning rate of $1\times 10^{-5}$, LoRA rank of 64, and use Adam optimization. Training on NeSyGeo-Caption used 1 epoch, while NeSyGeo-CoT used 2 epochs.
C.5 Diversity of Dataset
Diversity: Datasets synthesized by our NeSyGeo framework exhibit high diversity in text and visual features.
A critical challenge for automatic data synthesis methods is whether the dataset is sufficiently diverse to avoid quality degradation due to potential overfitting risks from inherent domain constraints. We employ t-SNE (van der Maaten & Hinton, 2008) dimensionality reduction for mapping in text space to evaluate the diversity across different methods. This analysis allows us to assess the diversity of the textual descriptions themselves. Given that in geometric problems, the text conditions depict specific visual elements, and the diversity observed in the text space also serves as a valuable indicator of the diversity in the corresponding visual diagrams. To ensure a fair comparison, we remove all prompts related to guiding large models, retaining only condition and question texts, and randomly sample 5k texts from each dataset. The results are illustrated in Figure 8. Our method and G-LLaVA (Gao et al., 2025) exhibit uniformly distributed features in the space, indicating low data overlap and high diversity. In contrast, R-CoT and MAVIS display varying degrees of clustered distribution, indicating more feature-similar samples.
Appendix D NeSyGeo-Test Benchmark
To probe the genuine cross-modal reasoning capabilities of current models, our empirical analysis is grounded in the NeSyGeo-test benchmark. This dataset was engineered to address the shortcomings of its predecessors. It comprises high-fidelity, programmatically rendered images and mitigates modality imbalance by ensuring crucial information is distributed across both the visual and textual inputs. A cornerstone of its design is the elimination of textual redundancies, a feature that compels models to ground their deductions in the visual context and prevents reliance on unimodal heuristics. Evaluation results on current mainstream open-source and closed-source MLLMs are shown in Table 9.
D.1 Quantitative Performance
A key observation is the benchmark’s considerable difficulty, which challenges even state-of-the-art MLLMs. This is particularly evident in the Hard difficulty tier, where the highest accuracy achieved by any closed-source model is merely 56.7%, while the leading open-source model, Qwen2.5-VL-32B-Instruct, scores only 43.3%. These results underscore the significant room for improvement in complex geometric reasoning. Furthermore, the effectiveness of our difficulty stratification is validated by the consistent performance degradation exhibited by most models as they progress from Easy to Medium and Hard questions. This trend confirms the rational design of our benchmark’s difficulty levels. Finally, the results reveal a clear performance gap between model types, with leading closed-source models like Claude-3.5-Sonnet (74.5%) consistently outperforming the best open-source models (55.8%), highlighting the current advantages of proprietary systems on such specialized tasks.
@@ D.1 Quantitative Performance
Qwen2.5-VL-7B-Instruct 7B 38.3 30.7 32.2 36.8 27.5 32.3 33.3
Qwen2.5-VL-32B-Instruct 7B 38.3 30.7 32.2 65.8 49.4 41.7 58.2
InternVL2.5-8B 8B 55.9 56.2 55.4 64.5 49.3 43.3 55.8
LLaVA-NeXT-7B 13B 15.7 15.6 16.0 15.8 16.0 15.2 15.8
LLaVA-NeXT-13B 7B 23.7 15.5 15.4 18.6 14.7 13.3 16.4
@@
D.2 Analysis of Failure Cases
Despite recent advancements, large multimodal models (LMMs) still exhibit several fundamental failure modes when tasked with complex geometric reasoning. Through our analysis, we identify and categorize three primary types of error:
Semantic Grounding Errors: This class of error occurs when the model fails to correctly associate visual annotations with their corresponding geometric entities. For instance, a model might misattribute the numerical value of a specific angle to an adjacent, unlabeled angle. While the model may occasionally identify such a contradiction through subsequent deductive steps, this process is unreliable. A more robust solution involves enhancing the model’s architecture to support iterative self-correction, allowing it to perform consistency checks across its reasoning chain and rectify initial grounding errors.
Deficiencies in Latent Property Inference: Models often overlook implicit geometric properties that are not explicitly annotated in the diagram, thereby increasing the perceived complexity of the problem. A common example is the failure to infer the congruence of sides in an isosceles triangle, which leads the model into a spurious exploration of the solution space through exhaustive, case-based analysis. This frequently results in wasted computational budget on incorrect deductive paths, sometimes causing premature termination due to exceeding context length limits. Future work should focus on improving the model’s perceptual acuity and multi-step reasoning capabilities, enabling it to prune invalid branches of logic efficiently, even when faced with incomplete information.
Vulnerability to Deductive Fallacies: The reasoning process can be derailed by irrelevant visual information or distractors within the image, causing the model to commit logical errors. This often results in the generation of self-contradictory statements or conclusions that are tangential to the problem’s solution path. To mitigate this, we propose the integration of a formal verification module. Such a neuro-symbolic validator would serve as a safeguard, programmatically ensuring that each step in the chain-of-thought is a logically sound consequence of the established premises and previously derived facts, thereby enforcing a more rigorous deductive process.
We present additional failure cases of Gemini-2.0-Flash in Figures 9 and 10, where the error types made by the model during the geometric reasoning task are highlighted with corresponding colors.
Figure 9: Failure cases of NeSyGeo-Test.
Figure 10: Failure cases of NeSyGeo-Test.
D.3 Future Direction
Addressing these challenges points toward several key research directions for advancing multimodal geometric problem-solving. The path forward likely involves a tighter integration of perceptual modules with symbolic reasoning engines, moving toward more capable architectures. Furthermore, progress will depend on the development of architectures that explicitly support stateful, long-horizon reasoning, allowing them to maintain and update a coherent ”mental model” of the geometric state. Finally, enabling models to achieve self-improvement through reinforcement learning loops, where rewards are provided by feedback from formal verifiers or successful problem completion, represents a crucial frontier for creating more robust and reliable geometric reasoning agents.
Appendix E Case Analysis
E.1 Comparison with other Popular Geometry Datasets
Figure 11: Comparison of NeSyGeo-CoT dataset with other Popular Geometry Datasets. Geometry-3K is a manually synthesized dataset, while the remaining approaches employ automatic generation techniques. Our dataset features clear, human-aesthetically pleasing images, high-quality step-by-step reasoning chains, symbolic form meta-information enabling subsequent image augmentation and mutation, and well-distributed conditional information between images and text.
E.2 More examples of our datasets
E.3 Case of Reasoner and Verifier
Figure 13: Case of reasoner and verifier.
Figure 14: Cases filtered out due to an inconsistency detected during the verification phase.
To provide a transparent illustration of our methodology, we present a concrete case study in Figure 13. This example details the step-by-step reverse search conducted by the Reasoner and the subsequent forward reasoning performed by the Verifier. Furthermore, to highlight the effectiveness of our validation mechanism, we also showcase an instance in Figure13that was filtered out due to an inconsistency detected during the verification phase.
Appendix F Use of LLMs
Within the Reasoner and Verifier modules of NeSyGeo, we employ LLMs to perform deductive geometric reasoning. This is actualized through a two-stage generation process, comprising a step-by-step reverse search and a reliable forward validation. The formidable reasoning capabilities and extensive pre-trained knowledge inherent in LLMs facilitate the efficient expansion of reasoning states, while simultaneously ensuring the fidelity and diversity of the resulting reasoning paths.
All technical contributions, including the design of experiments, comparison of MLLMs (e.g., Qwen and GPT-4o), and interpretation of results, were fully conceived and conducted by the authors. The authors take full responsibility for the correctness, originality, and integrity of all scientific content presented in the paper.
Appendix G Future Work
We intend to extend NeSyGeo to other multimodal domains, such as analytical geometry and visual question answering. This extensibility will be achieved by defining new domain-specific languages, corresponding synthesis rules within the symbolic space, and tailored conversion engines. Furthermore, we plan to develop an automated symbolic solver capable of conducting search and validation directly within the symbolic space. This would remove reliance on LLMs, potentially reducing generation costs and ensuring complete correctness of the datasets.
Appendix H Details of Domain-Specific Languages
H.1 Definition of Geo-DSL
@@
Geo-DSL adopts an entity–attributes-relations framework to define geometric elements in plane geometry, encompassing 13 types of points, 7 types of lines, 3 types of angles, and 14 types of shapes. Representative examples of symbolic statements and their corresponding natural language descriptions are illustrated in Figures 19, 16, 17, and 18. With a single statement, Geo-DSL uniquely specifies spatial elements, ensuring the accuracy of geometric synthesis while significantly facilitating parsing and transformation by our conversion engine. This language achieves comprehensive coverage of plane geometry, including numerical attributes such as lengths and angle measures, enabling precise and complete geometric representations. By streamlining definitions into concise statements, Geo-DSL reduces the complexity of symbolic processing, enhances the efficiency of the conversion engine, and supports seamless integration with neural synthesis pipelines. These advantages make Geo-DSL a robust and versatile solution for generating high-quality multimodal geometric reasoning data.
H.2 Definition of Actions in Symbolic Sequence Generation